answersLogoWhite

0

Lol basically yes :P

User Avatar

Wiki User

∙ 14y ago
User Avatar

What else can I help you with?

Continue Learning about Statistics

What is the product rule and the sum rule of probability?

Sum Rule: P(A) = \sum_{B} P(A,B) Product Rule: P(A , B) = P(A) P(B|A) or P(A, B)=P(B) P(A|B) [P(A|B) means probability of A given that B has occurred] P(A, B) = P(A) P(B) , if A and B are independent events.


How do you find P A given B?

P(A|B)= P(A n B) / P(B) P(A n B) = probability of both A and B happening to check for independence you see if P(A|B) = P(B)


Addition rule for probability of events A and B?

If they're disjoint events: P(A and B) = P(A) + P(B) Generally: P(A and B) = P(A) + P(B) - P(A|B)


Give the example of why probabilities of A given B and B given A are not same?

Let's try this example (best conceived of as a squared 2x2 table with sums to the side). The comma here is an AND logical operator. P(A, B) = 0.1 P(A, non-B) = 0.4 P(non-A, B) = 0.3 P(non-A, non-B) = 0.2 then P(A) and P(B) are obtained by summing on the different sides of the table: P(A) = P(A, B) + P(A, non-B) = 0.1 + 0.4 = 0.5 P(B) = P(A,B) + P(non-A, B) = 0.1 + 0.3 = 0.4 so P(A given B) = P (A, B) / P (B) = 0.1 / 0.4 = 0.25 also written P(A|B) P(B given A) = P (A,B) / P (A) = 0.1 / 0.5 = 0.2 The difference comes from the different negated events added to form the whole P(A) and P(B). If P(A, non-B) = P (B, non-A) then P(A) = P(B) and also P(A|B) = P(B|A).


Definition of additive law in probability?

This has to do with the union of events. If events A and B are in the set S, then the union of A and B is the set of outcomes in A or B. This means that either event A or event B, or both, can occur. P(A or B) = P(A) + P(B) - P(A and B) **P(A and B) is subtracted, since by taking P(A) + P(B), their intersection, P(A and B), has already been included. In other words, if you did not subtract it, you would be including their intersection twice. Draw a Venn Diagram to visualize. If A and B can only happen separately, i.e., they are independent events and thus P(A and B) = 0, then, P(A or B) = P(A) + P(B) - P(A and B) = P(A) + P(B) - 0 = P(A) + P(B)

Related Questions

List of congressman in the batangas province new?

As of my last update in October 2023, the congressmen representing Batangas province are Raneo Abu for the 1st District, Mario Vittorio "Marvey" A. Mariño for the 2nd District, and Ajuy "Atoy" B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B. P. B. A. B.


If A and B are independent events then are A and B' independent?

if P(A)>0 then P(B'|A)=1-P(B|A) so P(A intersect B')=P(A)P(B'|A)=P(A)[1-P(B|A)] =P(A)[1-P(B)] =P(A)P(B') the definition of independent events is if P(A intersect B')=P(A)P(B') that is the proof


What is the product rule and the sum rule of probability?

Sum Rule: P(A) = \sum_{B} P(A,B) Product Rule: P(A , B) = P(A) P(B|A) or P(A, B)=P(B) P(A|B) [P(A|B) means probability of A given that B has occurred] P(A, B) = P(A) P(B) , if A and B are independent events.


What is the formula for inclusive events?

The probability of inclusive events A or B occurring is given by P(A or B) = P(A) + P(B) - P(A and B), where P(A) and P(B) represent the probabilities of events A and B occurring, respectively.


How do you find P A given B?

P(A|B)= P(A n B) / P(B) P(A n B) = probability of both A and B happening to check for independence you see if P(A|B) = P(B)


Addition rule for probability of events A and B?

If they're disjoint events: P(A and B) = P(A) + P(B) Generally: P(A and B) = P(A) + P(B) - P(A|B)


Give the example of why probabilities of A given B and B given A are not same?

Let's try this example (best conceived of as a squared 2x2 table with sums to the side). The comma here is an AND logical operator. P(A, B) = 0.1 P(A, non-B) = 0.4 P(non-A, B) = 0.3 P(non-A, non-B) = 0.2 then P(A) and P(B) are obtained by summing on the different sides of the table: P(A) = P(A, B) + P(A, non-B) = 0.1 + 0.4 = 0.5 P(B) = P(A,B) + P(non-A, B) = 0.1 + 0.3 = 0.4 so P(A given B) = P (A, B) / P (B) = 0.1 / 0.4 = 0.25 also written P(A|B) P(B given A) = P (A,B) / P (A) = 0.1 / 0.5 = 0.2 The difference comes from the different negated events added to form the whole P(A) and P(B). If P(A, non-B) = P (B, non-A) then P(A) = P(B) and also P(A|B) = P(B|A).


Definition of additive law in probability?

This has to do with the union of events. If events A and B are in the set S, then the union of A and B is the set of outcomes in A or B. This means that either event A or event B, or both, can occur. P(A or B) = P(A) + P(B) - P(A and B) **P(A and B) is subtracted, since by taking P(A) + P(B), their intersection, P(A and B), has already been included. In other words, if you did not subtract it, you would be including their intersection twice. Draw a Venn Diagram to visualize. If A and B can only happen separately, i.e., they are independent events and thus P(A and B) = 0, then, P(A or B) = P(A) + P(B) - P(A and B) = P(A) + P(B) - 0 = P(A) + P(B)


What does a line reflection not preserve?

A reflection in a line l is a correspondence that pairs each point in the plane and not on the linewith point P' such that l is the perpendicular bisector of segment PP'. IF P is on l then P is paired with itself ... Under a reflection the image is laterally inverted. Thus reflection does NOT preserve orientation...


Formulas on Percentage Base and Rate?

P=B×RB=P÷RR=P÷B


How do you find the coordinates of a quadrilateral after reflecting it over the y axis?

If the coordinates of a point, before reflection, were (p, q) then after reflection, they will be (-p, q).


What is the reflection of point P-1 6 across the line y x?

To find the reflection of point P(-1, 6) across the line y = x, you swap the x and y coordinates of the point. Therefore, the reflection of P(-1, 6) is P'(6, -1).

Trending Questions
Which measure of central tendency would be most appropriate to use to describe the most common diagnosis among clients receiving treatment at an outpatient mental health clinc? What is the difference between payoff table and decision trees? What does the squiggly line mean when it starts at zero on a graph? Why does data presented in a line graphs? What are the mean median range and mode of 5436? What is the probability of obtaining exactly 4 tails from the 9 flips of the coin? How much should a rolex gmt master weigh? What is confidence associated with an interval estimate is called the? What is the probability of drawing a red ten or a black ace? Find the probability of this event if one of these cards is drawn at random. Selecting a U not replacing it and then selecting an M.? What is probability of drawing a diamond and a club from a deck of cards? What are the chances of getting pregnant if your partner has had a vasectomy? What is the relation between mean mode median in symmetric normal distribution? What percentage of data values of a normal distribution will fall within one standard deviation below the mean? What is the value of the digit 6 in 76.052? Is value of six sigma higher? A six sided die is rolled twice what is the probability of rolling either an even number on the first roll or a 1 on the second roll the answer is 7 over 12 but I can't figure out how to get there? What is 142.6473 rounded to the nearest hundredth? What is the probability of drawing Spade from a pack of cards and drawing an Ace from a pack of cards? What is the definition that fits the term mode?